OpenStudy (anonymous):

solve: c+s+r=13.5 6c+3s+r=42 r=2s

OpenStudy (shayaan_mustafa):

Hi krystilmrotte :) How are you?

OpenStudy (anonymous):

good how are you

OpenStudy (shayaan_mustafa):

So you have to find the values of c, s and r.

OpenStudy (shayaan_mustafa):

Me too fine dear.

OpenStudy (anonymous):

yup and im tottsaly confused on how to find them

OpenStudy (shayaan_mustafa):

First eliminate r from both upper two equations.

OpenStudy (shayaan_mustafa):

Put r=2s in, c+s+r=13.5 c+s+2s=13.5 c+3s=13.5 Put r=2s in 6c+3s+r=42 6c+3s+2s=42 6c+5s=42 Uptil now you have cleared?

OpenStudy (anonymous):

still confused....

OpenStudy (shayaan_mustafa):

Where ? What step you did not get?

OpenStudy (anonymous):

after the second "put r+2s"

OpenStudy (shayaan_mustafa):

I didn't written put r+2s anywhere. Look our first objective is to reduce the all three equations into the two equations. So therefore, when we say two equation then we mean that there must be two variables. Like in the question you have given with 3 equation so there must be 3 variables too. As we have namely, c, s and r.

OpenStudy (shayaan_mustafa):

Do you know this all?